Q2. Assume that the current exchange rate between $ and Russian Ruble is 30Rb/$. The Russian Central Bank announces that Ruble will not get weaker than 40Rb/$ during the next year. Assume that one-year Ruble interest rate is 10% and one-year $ interest rate is 5%. Do you find the Central Bank’s announcement credible? Explain. (Hint: What does Uncovered Interest Parity (UIP) suggest for the future spot rate?) (Show your calculations!)
Answer #1

Answer: Central Bank announcement is credible. Infact, future spot rate will not go above 31.42 Rb/$ as per UIP

A. According to Uncovered Interest parity (UIP),the difference in the interest rate between 2 nations is equal to expected change in exchange rates between those 2 countries' currencies.

In this case, The difference between 2 countries' interest rates = 10% - 5% = 5%. This means that Ruble is expected to depreciate by the same 5% (if uncovered interest parity (UIP) condition exists). So expected future spot rate would be 30*1.05 = 31.42Rb/$

More qualitatively, the below explanation would also hold good.

B. If a investor invests 100$ in US 1 year bond, then the return will be 100$ (1+5%) = $105

If a investor invests same 100$ in Russian 1 year bond, then the return will be 100$*current exchange rate*(1+10%) = 3300 ruble. If the future return is same as the investment in US (as per UIP), then $105 = Rb3300/expected future spot rate

So expected future spot rate would be 31.42 Rb/$